A car with good tires on a dry road can decelerate at about 5.0 m/s^2 when braking. Suppose a car is initially traveling at 55 mi/h. What is the stopping distance


1
Expert's answer
2021-09-14T09:35:44-0400

From the kinematic equation, the distance travelled by the car is the following (see https://www.khanacademy.org/science/physics/one-dimensional-motion/kinematic-formulas/a/what-are-the-kinematic-formulas):


"\Delta x = \dfrac{v_f^2 - v_i^2}{2a}"

where "v_f = 0m\/s" is the final speed (car stops at the end), "v_i = 55mi/h \approx 24.59 m/s" is the initial speed, and "a = -5m\/s^2" is the car's decceleration. Thus, obtain:


"\Delta x = \dfrac{0^2 - 24.59^2}{-2\cdot 5} \approx 60.47m"

Answer. 60.47 m.